Accelerator Update for March 21, 2008 - March 24, 2008
Friday March 21
The day shift began with the Tevatron (TeV) in access, with Operations monitoring the Recycler stash (233.4E10), with the Antiproton Source stacking (9.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) in access, and with NuMI and the SciBooNE / MiniBooNE (BNB) experiments taking beam.
Booster experts began work on a bulk power supply (CEP400) at 8:37 AM.
Interlock experts began testing the TeV sectors B, C, D, & E safety systems at 8:50 AM.
Operations prepared the Pbar pre-target and pre-vault areas for access at 9:40 AM.
A TeV expert reported at 10:22 AM that the B0 Quench Protection Monitor was operational.
Interlock experts began testing the TeV F-sector safety systems at 8:50 AM, with A-E sectors to follow.
Booster experts completed their work on CEP400 at 11:24 AM.
Booster experts took beam at 11:25 AM.
At 12:55 PM, a Linac RF station (LRF3) tripped off. It reset.
TeV experts reported at 1:03 PM that they had added coolant to the TeV sector F17 kickers.
At 1:43 PM, Operations reported that a Booster RF station (BRF9) had tripped off several times. An operator contacted an RF expert
CDF reported that they had completed their access. Operators search and secured the area at 3:42 PM.
Interlock experts completed testing the D0 safety system at 4:34 PM.
D0 came out of access at 5:37 PM.
Operations put the TeV through a dry squeeze at 7:58 PM.
At 8:19 PM, LRF3 tripped off. It reset.
At 8:23 PM, KRF4 tripped off. It reset.
Operations put the TeV through a wet squeeze at 9:12 PM.
At 9:19 PM, Operations reported that the Pbar PBKICK front end (FE) was not responding. Operations rebooted the FE.
At 11:53 PM, Operations reported a problem with a TeV quadrupole power supply (C:DQ0). An expert was called in.
Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.
Saturday March 22
The midnight shift began with the TeV shot setup on hold due to power supply work, with Operations monitoring the stash (310.7E10), with Pbar stacking (111.0E10), and with NuMI and BNB taking beam.
At 3:34 AM, the TeV power supply expert reported that C:DQ0 had a faulty transductor.
Operations reported at 4:56 AM that the C:DQ0 repair was completed.
At 4:58 AM, LRF3 tripped off. It reset.
At 5:32 AM, Operations reported that LRF3 had quite a few trips in the last hour. Operations called in a Linac expert.
Operations put the TeV through a dry squeeze at 5:42 AM.
At 6:21 AM, Operations reported that LRF3 continued to trip off.
Operations put the TeV through a wet squeeze at 6:39 AM.
Operations reported at 7:58 AM that the Recycler had stashed once during the midnight shift.
At 8:24 AM, a Linac expert took down LRF3 to make repairs. He had it back online in an hour.
Operations put the TeV into shot setup at 10:02 AM.
At 1:39 PM, the TeV sector B2 wet engine tripped off and would not reset. Cryo system experts at CHL opened the JT valve and the system was maintaining itself. An inspection showed that the engine brake had engaged and would not release.
CDF and D0 were both back to data taking by 2:32 PM. They had both turned off when Operations informed them that a wet engine had tripped off and that Operations might have needed to abort the store.
At 2:40 PM, KRF7 tripped off. It reset.
At 3:13 PM, BRF7 tripped off and caused the East Anode Power Supply (EAPS) to trip. They both reset.
Cryo system experts reported at 3:54 PM that they had repaired a brake sensor switch.
At 7:21 PM, a Booster sweeping magnet tripped off for the first of two times in about 20 minutes. The first trip was due to bad vacuum and reset. The second trip was due to a vacuum pump (IPMP02) tripping off. An expert said he’d come in and investigate.
Operations put Pbar into standby at 7:55 PM.
Operations reported at 8:02 PM that the Recycler had stashed six times during the last 12 hours.
Booster resumed sending beam at 8:17 PM. Pbar resumed stacking.
At 8:48 PM, a Booster expert began looking at problems with B:MP02 and its vacuum pump, IPMP02.
Sunday March 23
The midnight shift began with Operations monitoring store 6001 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
At 12:32 AM, MP02 tripped off. Operators disconnected the “B” cable from the IPMP02 power supply. Everything reset.
At 1:43 AM, KRF3 tripped off. It reset.
At 1:49 AM, B:IPMP02 tripped off and caused MP02 to also trip. Everything reset.
At 4:35 AM, the Pelletron tripped off. It reset.
Operations terminated store 6001 at 5:18 AM.
At 5:36 AM, the TeV suffered a sector B2 quench.
At 6:01 AM, BRF10 tripped off. It reset.
At 6:24 AM, Operations reported that the B2 wet engine tripped off. Cryo experts came in to rebuild the engine.
At 7:39 AM, LRF4 tripped off. It reset.
Operations reported at 8:00 AM that the Recycler had stashed five times during the last 12 hours.
Cryo experts reported at 9:41 AM that they had replaced four valves and replaced an “O” ring and did not need to do a complete rebuild.
Operations began the TeV turn on at 10:34 AM.
At 11 AM, MP02 tripped off. It reset.
Operations put the TeV into a dry squeeze at 11:28 AM.
Operations put the TeV into shot setup at 11:54 AM.
Operations established store 6003 at 1:38 PM with an initial luminosity of 245.02E30.
At 6:17 PM, BRF9 & 16 tripped off. They reset.
At 6:48 PM, the Linac Klystron vernier tripped off. It reset.
At 7:55 PM, The Linac beam permit tripped off due to a Spectrometer power supply trip (L:SPEC) in the 400MeV area. Operations contacted experts.
Operations established an energy-conserving time line at 8:34 PM.
A Linac expert set up to run beam to the straight through dump and on to Booster at 9:28 PM.
At 10:24 PM, a Booster expert, who was helping with the Linac problems, reported that the Spectrometer 480 V breaker was tripped. He reset the breaker and the power supply came back on, but it tripped when its output reached 81 amps. Another expert said that the spare power supply might need a slight modification.
Operations reported at 10:40 PM that Pbar and NuMI resumed taking beam, but at reduced rates. BNB was off.
At 11:05 PM, BRF1 tripped off. It reset.
Experts were looking for a spare supply at 11:12 PM that might work for L:SPEC.
BNB resumed taking beam at 11:59 PM.
Monday March 24
The midnight shift began with Operations monitoring store 6003 and the stash, with Pbar stacking, and with NuMI and BNB taking beam.
At 1:25 AM, Pbar and NuMI were turned off.
At 1:40 AM, LRF2 tripped off. It reset.
Operations reported that the first spare power supply for L:SPEC didn’t work. The second supply ran for a while, but beam positions were bad. Another Booster expert was called in to help.
Operations reported at 3:33 AM that experts had installed another L:SPEC power supply. Pbar, NuMI, and BNB resumed taking beam.
At 3:58 AM, the Linac Spectrometer power supply tripped off.
At 4:29 AM, experts worked at reconfiguring a different power supply for L:SPEC.
Linac resumed sending beam at 5:16 AM.
At 5:54 AM, L:SPEC tripped off. It reset.
Operations reported at 7:57AM that the Recycler had stashed two times during the last 12 hours.
The Plans for Monday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 03/24/2008 email Fermilab|
|Security, Privacy, Legal|